Find the slope. if it exists, of the line containing the pairs of points.
(-12.3,2.9) and (-11.9,2.7)
The slope m=____. Simplify your answer

Answer by checkley77(12844)

(-12.3,2.9) and (-11.9,2.7)
m=(y2-y1)/(x2-x1)
m=(2.7-2.9)/(-11.9+12.3)
m=-.2/.4
m=-1/2=.5 ans. for the slope.